# TiTok: Transfer Token-level Knowledge via Contrastive Excess to Transplant LoRA

### v1

@@ Abstract
Large Language Models (LLMs) are widely applied in real world scenarios, but fine-tuning them comes with significant computational and storage costs. Parameter-Efficient Fine-Tuning (PEFT) methods such as LoRA mitigate these costs, but the adapted parameters are dependent on the base model and cannot be transferred across different backbones. One way to address this issue is through knowledge distillation, but its effectiveness inherently depends on training data. Recent work such as TransLoRA avoids this by generating synthetic data, but this adds complexity because it requires training an additional discriminator model. In this paper, we propose TiTok [^1] Code will be released upon acceptance: https://github.com/NaughtyMaltiz16/TiTok, a new framework that enables effective LoRA Transplantation through Token-level knowledge transfer. Specifically, TiTok captures task-relevant information through a contrastive excess between a source model with and without LoRA. This excess highlights informative tokens and enables selective filtering of synthetic data, all without additional models or overhead. Through experiments on three benchmarks across multiple transfer settings, our experiments show that the proposed method is consistently effective, achieving average performance gains of + 4–8% compared to baselines overall.
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Vaswani et al., 2023), have made significant progress in many real-world applications, including chatbots (OpenAI et al., 2024), search engines (Xiong et al., 2024), and coding assistants (Rozière et al., 2024). While fine-tuning LLMs has been demonstrated to be a promising way to improve performance on downstream tasks, it incurs substantial computational and storage costs. Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019) methods such as LoRA (Hu et al., 2021) alleviate this burden by updating only a small subset of parameters while keeping the base model frozen. However, PEFT’s adapted parameters are dependent to the base model and cannot be transferred across different models. This limitation is increasingly critical given the rapid release of new LLMs and the growing diversity of available models.
@@
One considerable approach to mitigate this limitation is through Knowledge distillation (KD) (Hinton et al., 2015; Azimi et al., 2024), which transfers the knowledge embedded in a source model’s PEFT adapters to a target model with new PEFT adapters by aligning the target’s output distributions with those of the source. However, KD is inherently data-dependent, typically requiring access to training data from target downstream tasks (Nayak et al., 2019; Liu et al., 2024), which is often unavailable or costly to obtain. To address this limitation, TransLoRA (Wang et al., 2024) has recently proposed using synthetic data by leveraging the data synthesis capabilities of recent LLMs (Wang et al., 2023; Kim et al., 2025). This approach enables the target model to acquire domain knowledge without direct access to the original dataset. Nevertheless, TransLoRA requires training an additional discriminator model to filter low-quality synthetic data, which inevitably introduces extra complexity and computational overhead. Furthermore, it primarily emphasizes the role of synthetic data, paying less attention to how the knowledge transfer process itself should be designed.
@@
Figure 1: Overview of TiTok: Transplantation through Token-level knowledge transfer. Starting from a small set of seed prompts, the source expert model (source base model + LoRA) generates synthetic data. A contrastive excess filtering mechanism then compares the expert against its base backbone to compute token-level excess scores. Using these scores, TiTok first performs sample filtering and subsequently token selection, retaining only the most informative samples and tokens. When tokenizers differ, masks are aligned prior to training. The resulting filtered data is finally used to train a new LoRA on the target backbone, enabling efficient knowledge transfer.
@@
Contribution. In this paper, we propose a new framework that enables effective LoRA Transplantation through Token-level knowledge transfer (TiTok). Our high-level idea is to selectively convey task-relevant information from the source model’s LoRA by using token-level signals to guide the transfer process, rather than relying on the entire token sequence. We specifically capture this information through a concept we introduce as contrastive excess, obtained by comparing predictions from a source model with LoRA and the same model without LoRA. Intuitively, this contrastive excess highlights tokens that contain important task knowledge. This excess signal is further utilized to filter the synthetic data we generate for training, thereby enabling selective learning on samples that contain richer information. Unlike TransLoRA, which requires training and storing an additional discriminator model for filtering, TiTok requires neither extra models nor additional training overhead. Moreover, we design an effective mechanism to resolve tokenizer mismatches between source and target models, which enhances robustness and applicability of our framework.
@@
We demonstrate the effectiveness of TiTok by conducting extensive experiments on three widely used benchmarks, which cover both reasoning (Big Bench Hard (Suzgun et al., 2022) and MMLU (Hendrycks et al., 2021)) and personalization tasks (LaMP (Salemi et al., 2024)). In particular, TiTok improves the performance by +7.96% over the vanilla target model, +6.0% over KD, and +4.4% over TransLoRA, when averaged across all tasks and transfer settings. We also explored a variety of transfer settings, including transfers within the same model, across different model families and sizes, and even across different model versions. In every case, our approach achieves consistent improvements. Interestingly, TiTok remains effective even when applied to external data originating from tasks different from the target task, highlighting both robustness and general applicability. Overall, these empirical results highlight TiTok as a methodologically simple yet powerful paradigm for efficiently transferring LoRA knowledge across models in diverse scenarios.
@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019; Li & Liang, 2021) has emerged as both a practical and popular alternative to full model fine-tuning. By requiring updates to only a small portion of parameters, it enables efficient adaptation, with LoRA (Low-Rank Adaptation) (Hu et al., 2021; Dettmers et al., 2023) standing out as one of the most widely adopted methods. However, a fundamental limitation is that LoRA adapters are tied to the frozen backbone they were trained on, making them difficult to transfer to other base models. To address this issue, recent studies such as TransLoRA (Wang et al., 2024) have attempted to transplant the knowledge of LoRA adapters across models by generating synthetic data (Wang et al., 2023). While effective to some extent, this approach requires an additional discriminator model to filter high-quality synthetic data, resulting in a relatively heavy pipeline. In parallel, Knowledge Distillation (KD) (Hinton et al., 2015; Azimi et al., 2024) has been widely explored as another means of knowledge transfer; however, traditional KD typically operates within a teacher–student framework at the logit or sequence level and requires access to training data close to the original, in order to use the teacher’s distribution for supervising the student. In contrast, our approach focuses on token-level selective transfer, offering a more fine-grained yet lightweight alternative that enables efficient transplantation of LoRA adapters across diverse models for deployment.
@@ 3 TiTok: Transplanting LoRA through Token-Level Knowledge
In this section, we introduce TiTok, a framework for LoRA Transplantation through Token-level knowledge transfer (Fig. 1). The core idea of TiTok is to transfer the knowledge from a source model’s LoRA adapter into a target model’s LoRA adapter by training specifically on the informative tokens within synthetic data. Specifically, the framework consists of three components: (1) Synthetic data generation (Sec. 3.1), where the source expert model produces query–label pairs for target task; (2) Excess score computation (Sec. 3.2), which calculates token-level importance using source model; and (3) Target model training with filtering (Sec. 3.3), which trains target model with newly initialized LoRA adapter on top-ranked samples and tokens. In addition, we propose Excess score alignment (Sec. 3.4), an algortihm designed to apply TiTok even when tokenizers of the source and target models differ. The overall algorithm of TiTok is presented in Alg. 1.
@@ 3.1 Synthetic data generation via LLM prompting with few-shot data
Let $\mathcal{M}_{s}$ denote the source backbone LLM and $\mathcal{A}_{s}$ its LoRA adapter on target downstream task, which forms the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$. The target model, whose LoRA adapter $\mathcal{A}_{t}$ will be trained, is denoted by $\mathcal{M}_{t}$. Then, TiTok first constructs a synthetic dataset $\mathcal{D}_{\text{s}}$, similar to the idea of TransLoRA (Wang et al., 2024). This usage of synthetic data allows us to avoid keeping the whole original dataset for the downstream task, and simultaneously let $\mathcal{A}_{t}$ learn the knowledge encoded in the source adapter $\mathcal{A}_{s}$. Unlike TransLoRA’s approach of using the untuned target model $\mathcal{M}_{t}$ to generate synthetic data, we use the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$ to synthesize data (see empirical comparison in Fig. 2). Concretely, $\mathcal{M}_{s}+\mathcal{A}_{s}$ synthesizes both the query and the label within a prompting-based data synthesis framework (Wang et al., 2023) (see details in Appendix H); given a few-shot data of downstream task, it first generates a query q, and then produces the corresponding label y conditioned on q. To encourage diversity, we apply ROUGE-L filtering together with deduplication to all tasks, except for exceptional cases where such filtering is infeasible (more details are in Appendix G). Consequently, the resulting synthetic dataset consists of query–label pairs:
@@
3.2 Contrastive excess score from source model with LoRA adapters
As described in Sec. 3.1, TiTok relies on synthetic data, but synthetic data is often prone to imperfections (Chen et al., 2024); thus, sophisticated filtering is essential to retain high-quality informative samples. TransLoRA (Wang et al., 2024) tackles this challenge with a separate discriminator to filter useful queries, but this introduces the extra burden of training and maintaining an extra model. In contrast, we propose a lightweight alternative that only utilizes the already trained source model. Specifically, we use source model and its LoRA adapter to perform two complementary roles: (1) the amateur role ($\mathcal{M}_{s}$) and (2) the expert role ($\mathcal{M}_{s}+\mathcal{A}_{s}$). Then, the difference between the two roles provides an implicit supervision signal where the task information is encoded.
@@ 3.4 Excess score alignment across different tokenizers
In cases of transfer between models with different tokenizers, a direct mapping of token-level signals is not possible, given that the source and target models may segment text differently. To address this, we introduce a simple yet robust tokenizer alignment algorithm that propagates token masks (Eq. 6) from the source token sequence $\mathbf{y}^{(s)}$ to the target token sequence $\mathbf{y}^{(t)}$. The algorithm first aligns token sequences using dual pointers that incrementally decode and match text spans. Masks are then propagated using the following four rules: (1) direct copy for one-to-one mappings, (2) replication for one-to-many, (3) averaging for many-to-one, and (4) averaging with replication for many-to-many. Finally, a top-$k\%$ selection step retains the most confident target tokens. This process ensures consistent supervision across tokenizers, enabling reliable transfer even when models tokenize text differently. The conceptual illustration of this procedure is presented in Fig.4.
@@ 4.1 Experimental setups
To demonstrate the effectiveness of TiTok, we compare it against three baselines: (i) Vanilla, the target base model without any fine-tuning; (ii) KD(+MinED), knowledge distillation (KD) from source expert model (Hinton et al., 2015). When the source and target models use different tokenizers, the original KD is not applicable. In this case, we use Minimum-Edit-Distance (MinED) tokenizer alignment (Wan et al., 2024), which aligns both token sequences and vocabulary distributions via dynamic-programming sequence alignment for near matches (e.g., “gets” $\leftrightarrow$ “get”) and by mapping probability mass to nearest edit-distance neighbors (e.g., “immediately” $\leftrightarrow$ “immediate”). We use the synthesized data by TransLoRA as the training data for KD and MinED ; and (iii) TransLoRA (Wang et al., 2024), a prior method in which the vanilla target model synthesizes the queries, while the source model with its LoRA adapter generates the corresponding synthetic labels. A discriminator is then employed to filter the synthetic data for training the target LoRA adapter.
@@
Following prior work in TransLoRA, we first conduct experiments on two representative benchmarks: (1) Big-Bench Hard (BBH) (Suzgun et al., 2022) and (2) Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021). BBH consists of 27 challenging reasoning tasks structured as multiple choice or short answer questions, designed to test compositional generalization and advanced problem solving abilities of a Language Model. Meanwhile, MMLU covers 57 tasks across diverse academic subjects, presented in multiple choice format to evaluate broad knowledge and reasoning skills of a model. Since both benchmarks only provide test sets, we split the data into 90% for training the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$ and the remaining 10% for evaluation.
To extend our approach to personalization and text generation, we additionally conduct experiments on LaMP benchmark (Salemi et al., 2024), focusing exclusively on its generation tasks. In particular, we experiment with (3) News Headline Generation (LaMP 4) and (4) Scholarly Title Generation (LaMP 5), as they are the only text generation tasks that are both accessible and reliably evaluable. The remaining LaMP tasks are excluded, given that LaMP 1-3 are discriminative, and LaMP 6-7 lack gold labels. For LaMP tasks, the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$ is trained on data from the 30 users with the longest activity histories. From each user, we use 200 data points for training and 50 data points for validation, a design choice intended to conduct a more rigorous and robust evaluation. In total, each LaMP task contains 6,000 training examples and 1,500 evaluation examples.
@@
For training of both source and target models, we use a learning rate of $5\times 10^{-5}$, train for 2 epochs with a batch size of 4, and apply LoRA with rank $r=8$, scaling factor $\alpha=8$, and dropout $0.05$. Optimization is performed using AdamW with weight decay $1\times 10^{-2}$, together with a linear learning rate schedule and a warmup ratio of $0.1$. For synthetic data generation, we provide five samples from the original training data as few-shot exemplars, and apply top-p sampling (Holtzman et al., 2020) to generate both queries and labels, with sampling hyperparameters tuned individually for each task. To further encourage diversity, we apply ROUGE-L filtering with a threshold of $0.7$ and deduplication to remove redundant queries, following Wang et al. (2023). For tasks where ROUGE-based filtering is infeasible, we apply only deduplication (see Appendix G). For the initial synthetic pool, we generate $2M$ synthetic samples and, after filtering with contrastive excess scores, retain the top $M$, where $M$ equals the source training set size (see Appendix F). For token selection, the selection ratio $k\%$ is fixed at $70\%$ across all tasks and transfer settings, except for the Llama3 3B $\rightarrow$ Llama3 8B transfer setting, where we find that $k\%=30\%$ consistently yields the best performance. For inference, we adopt greedy decoding to ensure deterministic evaluation.

### v2

@@ Abstract
Large Language Models (LLMs) are widely applied in real world scenarios, yet fine-tuning them comes with significant computational and storage costs. Parameter-Efficient Fine-Tuning (PEFT) methods such as LoRA mitigate these costs; however, the adapted parameters are dependent on the base model and cannot be transferred across different backbones. One way to address this issue is through knowledge distillation, but its effectiveness inherently depends on training data. Recent work such as TransLoRA avoids this by generating synthetic data; nevertheless, this adds complexity since it requires training an additional discriminator model. In this paper, we propose TiTok, a new framework that enables effective LoRA Transplantation through Token-level knowledge transfer. Specifically, TiTok captures task-relevant information through a token-wise contrastive excess between a source model with and without LoRA. This excess highlights informative tokens and enables selective filtering of synthetic data, all without additional models or overhead. Through experiments on three benchmarks across multiple transfer settings, we demonstrate that TiTok is consistently effective, achieving average performance gains of + 4–10% compared to baselines overall. [^1] Code available at https://github.com/NaughtyMaltiz16/TiTok
@@ 1 Introduction
Large Language Models (LLMs) (Brown et al., 2020; Vaswani et al., 2023) have made significant progress in many real-world applications, including chatbots (OpenAI et al., 2024), search engines (Xiong et al., 2024), and coding assistants (rozière2024codellamaopenfoundation). While fine-tuning LLMs has been demonstrated to be a promising way to improve performance on downstream tasks, it incurs substantial computational and storage costs. To alleviate this, Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019) methods such as LoRA (Hu et al., 2021) update only a small subset of parameters while keeping the base model frozen. However, PEFT’s adapted parameters are dependent to the base model and thus cannot be transferred across different models. This limitation is increasingly critical given the rapid release of new LLMs and the growing diversity of available models.
@@
One potential approach to mitigate this limitation is Knowledge distillation (KD) (Hinton et al., 2015; Azimi et al., 2024), which transfers the knowledge embedded in a source model’s PEFT adapters to a target model with new PEFT adapters by aligning the target’s output distributions with those of the source. However, KD is inherently data-dependent, typically requiring access to training data from target downstream tasks (Nayak et al., 2019; Liu et al., 2024), which is often unavailable or costly to obtain. To address this limitation, TransLoRA (Wang et al., 2024) has recently proposed to use synthetic data, leveraging the data synthesis capabilities of recent LLMs (Wang et al., 2023; Kim et al., 2025). This approach enables the target model to acquire domain knowledge without direct access to the original dataset. Nevertheless, TransLoRA requires training an additional discriminator model to filter low-quality synthetic data, which inevitably introduces extra complexity and computational overhead. Furthermore, it primarily emphasizes the role of synthetic data, paying less attention to how the knowledge transfer process itself should be designed.
@@
Figure 1: Overview of TiTok: Transplantation through Token-level knowledge transfer. Starting from a small set of seed prompts, the source expert model (source base model + LoRA) generates synthetic data. A token-wise contrastive excess filtering mechanism then compares the expert against its base backbone to compute token-level excess scores. Using these scores, TiTok first performs sample filtering and subsequently token selection, retaining only the most informative samples and tokens. When tokenizers differ, masks are aligned prior to training. The resulting filtered data is finally used to train a new LoRA on the target backbone, enabling efficient knowledge transfer.
@@
Contribution. In this paper, we propose a new framework that enables effective LoRA Transplantation through Token-level knowledge transfer (TiTok). Our high-level idea is to selectively convey task-relevant information from the source model’s LoRA by using token-level signals to guide the transfer process, rather than relying on the entire token sequence. We specifically capture this information through a concept we introduce as token-wise contrastive excess, which is obtained by comparing predictions from a source model with LoRA and the same model without LoRA. Intuitively, this token-wise contrastive excess highlights tokens that contain important task knowledge. This excess signal is further utilized to filter the generated synthetic data for training, thereby enabling selective learning on samples that contain richer information. Unlike TransLoRA, which requires training an additional discriminator model for filtering, TiTok requires neither extra models nor additional training overhead. Moreover, we design an effective mechanism to resolve tokenizer mismatches between source and target models, enhancing robustness and applicability.
@@
We demonstrate the effectiveness of TiTok by conducting extensive experiments on three widely used benchmarks, which cover both reasoning (Big Bench Hard (Suzgun et al., 2022) and MMLU (Hendrycks et al., 2021)) and personalization tasks (LaMP (Salemi et al., 2024)). In particular, TiTok improves the performance by +9.94% over the vanilla target model, +8.5% over KD, and +4.4% over TransLoRA, when averaged across all tasks and transfer settings. We also explored a variety of transfer settings, including transfers within the same model, across different model families and sizes, and even across different model versions. In every case, our approach achieves consistent improvements. Interestingly, TiTok remains effective even when applied to external data originating from tasks different from the target task, highlighting both robustness and general applicability. Overall, these empirical results highlight TiTok as a methodologically simple yet powerful paradigm for efficiently transferring LoRA knowledge across models in diverse scenarios.
@@ 2 Related Works
Parameter-Efficient Fine-Tuning (PEFT) (Houlsby et al., 2019; Li and Liang, 2021) has emerged as both a practical and popular alternative to full model fine-tuning. By requiring updates to only a small portion of parameters, it enables efficient adaptation, with LoRA (Low-Rank Adaptation) (Hu et al., 2021; Dettmers et al., 2023) standing out as one of the most widely adopted methods. However, a fundamental limitation is that LoRA adapters are tied to the frozen backbone they were trained on, making them difficult to transfer to other base models. To address this issue, recent studies such as TransLoRA (Wang et al., 2024) have attempted to transplant the knowledge of LoRA adapters across models by generating synthetic data (Wang et al., 2023). While effective to some extent, this approach requires an additional discriminator model to filter high-quality synthetic data, resulting in a relatively heavy pipeline. In parallel, Knowledge Distillation (KD) (Hinton et al., 2015; Azimi et al., 2024) has been widely explored as another means of knowledge transfer; however, traditional KD typically operates within a teacher–student framework at the logit or sequence level and requires access to training data close to the original, in order to use the teacher’s distribution for supervising the student. In contrast, our approach focuses on token-level selective transfer, offering a more fine-grained yet lightweight alternative that enables efficient transplantation of LoRA adapters across diverse models for deployment.
@@ 3 TiTok: Transplanting LoRA through Token-Level Knowledge
In this section, we introduce TiTok, a framework for LoRA Transplantation through Token-level knowledge transfer (Fig. 1). The core idea of TiTok is to transfer the knowledge from a source model’s LoRA adapter into a target model’s LoRA adapter by training selectively on the informative tokens within synthetic data. Specifically, the framework consists of three components: (1) Synthetic data generation (Sec. 3.1), where the source expert model produces query–label pairs for target task; (2) Excess score computation (Sec. 3.2), which calculates token-level importance using the source model; and (3) Target model training with filtering (Sec. 3.3), which trains the target model with newly initialized LoRA adapter on top-ranked samples and tokens. In addition, we propose Excess score alignment (Sec. 3.4), an algortihm designed to apply TiTok even when tokenizers of the source and target models differ. The overall algorithm of TiTok is presented in Alg. 1.
@@ 3.1 Synthetic data generation via LLM prompting with few-shot data
Let $\mathcal{M}_{s}$ denote the source backbone LLM and $\mathcal{A}_{s}$ its LoRA adapter on target downstream task, which forms the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$. The target model, whose LoRA adapter $\mathcal{A}_{t}$ will be trained, is denoted by $\mathcal{M}_{t}$. Then, TiTok first constructs a synthetic dataset $\mathcal{D}_{\text{s}}$, similar to the idea of TransLoRA (Wang et al., 2024). This usage of synthetic data allows us to avoid keeping the whole original dataset for the downstream task, and simultaneously let $\mathcal{A}_{t}$ learn the knowledge encoded in the source adapter $\mathcal{A}_{s}$. Unlike TransLoRA’s approach of using the untuned target model $\mathcal{M}_{t}$ to generate synthetic data, we use the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$ to synthesize data (see empirical comparison in Fig. 3). Concretely, $\mathcal{M}_{s}+\mathcal{A}_{s}$ synthesizes both the query and the label within a prompting-based data synthesis framework (Wang et al., 2023) (see details in Appendix P); given a few-shot data of downstream task, it first generates a query q, and then produces the corresponding label y conditioned on q. To encourage diversity, we apply ROUGE-L filtering together with deduplication to all tasks, except for exceptional cases where such filtering is infeasible (more details are in Appendix I). Consequently, the resulting synthetic dataset consists of query–label pairs:
@@
3.2 Token-wise contrastive excess score from source model with LoRA
As described in Sec. 3.1, TiTok relies on synthetic data, but synthetic data is often prone to imperfections (Chen et al., 2024b); thus, sophisticated filtering is essential to retain high-quality informative samples. TransLoRA (Wang et al., 2024) tackles this challenge with a separate discriminator to filter useful queries, but this introduces the extra burden of training and maintaining an extra model. In contrast, we propose a lightweight alternative that only utilizes the already trained source model. Specifically, we use source model and its LoRA adapter to perform two complementary roles: (1) the amateur role ($\mathcal{M}_{s}$) and (2) the expert role ($\mathcal{M}_{s}+\mathcal{A}_{s}$). Then, the difference between the two roles provides an implicit supervision signal where the task information is encoded.
@@ 3.2 Token-wise contrastive excess score from source model with LoRA
While the excess score is motivated from an intuitive contrast between the backbone and its LoRA-enhanced counterpart, it is also grounded in principled statistical theory. Theoretically, the proposed token-wise contrastive excess score is connected to a token-level log-likelihood ratio (LLR) between the source expert (backbone + LoRA) and the backbone. LLR is a robust metric widely established in statistical testing (Li and Babu, 2019). For example, by the Neyman–Pearson lemma (Neyman and Pearson, 1933), the likelihood ratio is known to be the optimal statistic for identifying differences between two models. Therefore, high-LLR tokens are exactly the regions where the adapter changes the predictive distribution, i.e., where task knowledge is injected. In addition, from the perspective of standard knowledge distillation, tokens with nearly zero LLR provide no additional teacher knowledge and offer no benefit for transfer. In contrast, high-LLR tokens correspond to maximal teacher–student divergence, larger gradients, and higher information contribution to the student. Therefore, emphasizing tokens with high LLR in particular naturally focuses training on the precise positions where transferable knowledge is concentrated.
@@ 3.4 Excess score alignment across different tokenizers
In cases of transfer between models with different tokenizers, a direct mapping of token-level signals is not possible, given that the source and target models may segment texts differently. To address this, we introduce a simple yet robust tokenizer alignment algorithm that propagates token masks (Eq. 6) from the source token sequence $\mathbf{y}^{(s)}$ to the target token sequence $\mathbf{y}^{(t)}$. The algorithm first aligns token sequences using dual pointers that incrementally decode and match text spans. Masks are then propagated using the following four rules: (1) direct copy for one-to-one mappings, (2) replication for one-to-many, (3) averaging for many-to-one, and (4) averaging with replication for many-to-many. Finally, a top-$k\%$ selection step retains the most confident target tokens. This process ensures consistent supervision across tokenizers, enabling reliable transfer even when models tokenize text differently. The conceptual illustration of this procedure is presented in Fig.4.
@@ 4.1 Experimental setups
To demonstrate the effectiveness of TiTok, we compare it against three baselines: (i) Vanilla, the target base model without any fine-tuning; (ii) KD(+MinED), knowledge distillation (KD) from source expert model (Hinton et al., 2015). When the source and target models use different tokenizers, the original KD is not applicable. In this case, we use Minimum-Edit-Distance (MinED) tokenizer alignment (Wan et al., 2024), which aligns both token sequences and vocabulary distributions via dynamic-programming sequence alignment for near matches (e.g., “gets” $\leftrightarrow$ “get”) and by mapping probability mass to nearest edit-distance neighbors (e.g., “immediately” $\leftrightarrow$ “immediate”). We use the synthesized data by TransLoRA as the training data for KD (+MinED) ; and (iii) TransLoRA (Wang et al., 2024), a prior method in which the vanilla target model synthesizes the queries, while the source model with its LoRA adapter generates the corresponding synthetic labels. A discriminator is then employed to filter the synthetic data for training the target LoRA adapter.
@@
Following prior work in TransLoRA, we first conduct experiments on two representative benchmarks: (1) Big-Bench Hard (BBH) (Suzgun et al., 2022) and (2) Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021). BBH consists of 27 challenging reasoning tasks structured as multiple choice or short answer questions, designed to test compositional generalization and advanced problem solving abilities of a Language Model. Meanwhile, MMLU covers 57 tasks across diverse academic subjects, presented in multiple choice format to evaluate broad knowledge and reasoning skills of a model. As both benchmarks provide only test sets, we construct 90%/10% train-evaluation splits. For BBH, we use the full dataset, whereas for MMLU, we randomly sample 100 instances of the dataset for each task before applying the 90%/10% split.
To extend our approach to personalization and text generation, we additionally conduct experiments on LaMP benchmark (Salemi et al., 2024), focusing exclusively on its generation tasks. In particular, we experiment with (3) News Headline Generation (LaMP 4) and (4) Scholarly Title Generation (LaMP 5), as they are the only text generation tasks that are both accessible and reliably evaluable. The remaining LaMP tasks are excluded, given that LaMP 1,2,3 are discriminative, and LaMP 6,7 lack gold labels. For LaMP tasks, the source expert model $\mathcal{M}_{s}+\mathcal{A}_{s}$ is trained on data from the 30 users with the longest activity histories. From each user, we use 200 data points for training and 50 data points for validation, a design choice intended to conduct a more rigorous and robust evaluation. In total, each LaMP task contains 6,000 training examples and 1,500 evaluation examples.
@@
When training both the source and target models, we use a learning rate of $5\times 10^{-5}$, train for 2 epochs with a batch size of 4, and apply LoRA with rank $r=8$, scaling factor $\alpha=8$, and dropout $0.05$. Optimization is performed using AdamW with weight decay $1\times 10^{-2}$, together with a linear learning rate schedule and a warmup ratio of $0.1$. For synthetic data generation, we provide five samples from the original training data as few-shot exemplars, and apply top-p sampling (Holtzman et al., 2020) to generate both queries and labels, with sampling hyperparameters tuned individually for each task. To further encourage diversity, we apply ROUGE-L filtering with a threshold of $0.7$ and deduplication to remove redundant queries, following Wang et al. (2023). For tasks where ROUGE-based filtering is infeasible, we apply only deduplication (see Appendix I). For the initial synthetic pool, we generate $2M$ synthetic samples and, after filtering with token-wise contrastive excess scores, retain the top $M$, where $M$ equals the source training set size (see Appendix H). For token selection, the selection ratio $k\%$ is fixed at $70\%$ across all tasks and transfer settings, except for the Llama3 3B $\rightarrow$ Llama3 8B transfer setting, where we find that $k\%=30\%$ consistently yields the best performance. When inferencing during the evaluation stage, we adopt greedy decoding to ensure fully deterministic and reproducible inference results.